How many different committees of 4 members can be formed from a group with 7 seniors and 6 juniors if they are 3 seniors and 1 junior?
Answer by Edwin McCravy(20054)

7 seniors choose 3.  That's C(7,3), or 7C3, which is 35 ways to choose the 3 seniors.

For each of those 35 ways to choose the 3 seniors, there are 6 ways, 6C1, to choose the 1 junior.

Answer = C(7,3)×C(6,1) = 35×6 = 210 ways.

Edwin